Pranks, Inc. is a manufacturer of joke and novelty products for perpetrators of practical jokes. The corporation has paid several cash dividends throughout Year 6, the current year. It is also declaring a stock dividend to its stockholders as the calendar year-end approaches. You’ve been brought in as a consultant to assist with this process, and also to help determine whether some missing information can be determined before the distribution of the stock dividend is made. The company has two classes of stock: common stock and cumulative preferred stock.

You’ve been able to retrieve the following information so far:

Number of common shares authorized 800,000
Number of common shares issued 650,000
Par value of common shares $20
Par value of cumulative preferred shares $30
Paid-in capital in excess of par-common stock $7,000,000
Paid-in capital in excess of par-preferred stock $0
Total retained earnings before the stock dividend is declared $33,500,000
No treasury shares have been reissued.

Total Cash

Preferred Dividends

Common Dividends

Year

Dividends

Total

Per Share

Total

Per Share

Year 1 20,000 20,000 0.20 0 0.00
Year 2 36,000 36,000 0.36 0 0.00
Year 3 88,000 34,000 0.34 54,000 0.09
Year 4 120,000 30,000 0.30 90,000 0.15
Year 5 138,000 30,000 0.30 108,000 0.18
Year 6 210,000 30,000 0.30 180,000 0.30

Cash Dividends:

The accounting manager for the company prepared the schedule of cash dividends paid from Year 1 to Year 6 on the Pranks, Inc. panel. However, one of the reasons for Pranks, Inc.’s missing information is that the manager is away on vacation and is unreachable by phone, because he is backpacking on a remote island that does not have cell phone reception. Management would like you to determine some information from the data you’ve collected regarding its outstanding stock.

Fill in the following answers.

How many shares of common stock are outstanding?   
How many shares of preferred stock are outstanding?
What is the preferred dividend as a percent of par? %

Stock Dividend:

The company declared a 2% common stock dividend on December 1, and would like you to compute the following pieces of missing information. The market value of the common shares is $26.00 on December 1, and is $32.00 on the actual distribution date of the stock, December 31.

Fill in the missing information in the following table, using the information given and your work on the other panels. All “before” items are before the stock dividend was declared. All “after” items are after the stock dividend was declared and closing entries were recorded at the end of the year.

Total paid-in capital before the stock dividend   
Total retained earnings before the stock dividend
Total stockholders’ equity before the stock dividend
Total paid-in capital after the stock dividend
Total retained earnings after the stock dividend
Total stockholders’ equity after the stock dividend
Answer #1
Outstanding stock:
Common dividend is declared on common stock outstanding
Year 6,Total common dividend=$ 180000
Year 6,Common dividend per share=0.30
Common stock outstanding=180000/0.30=600000 shares
Preferred stock outstanding=Total preferred dividends in Year 6/Preferred dividends per share in year 6=30000/0.30=100000 shares
Preferred dividend as a % of par=Preferred dividends per share in year 6/par value of cumulative preferred shares=0.30/30=.01=1%
Stock dividend declared
When stock dividend is less than 20-25%, it is considered as small stock dividend
Stock dividend is declared on common stock outstanding
Common stock issued=600000*2%=12000 shares
Journal entry:
Date Debit Credit
Dec 1. Retained earnings (600000*2%*26) 312000
Common stock dividend distributable (600000*2%*20) 240000
Paid-in capital in excess of par-common stock 600000*2%*(26-20) 72000
(Stock dividend declared)
Dec 30. Common stock dividend distributable 240000
Common stock 240000
(Stock dividend issued)
Total paid-in capital before stock dividend:
$
Common stock (650000*20) 13000000
Preferred stock (100000*30) 3000000
Paid-in capital in excess of par-common stock 7000000
Paid-in capital in excess of par-preferred stock 0
Total paid-in capital 23000000
Note: the market value of common shares on Date of declaration is relevant and on the date of actual distribution is not relevant
Total retained earnings before the stock dividend=$ 33500000
Total stockholder's equity before the stock dividend=Total paid-in capital+Total retained earnings-Treasury stock
Number of shares in treasury stock=Number of common shares issued-Number of common shares outstanding=650000-600000=50000
Let's assume that treasury stock acquired at par value
Treasury stock in $=500000*20=$ 10000000
Total stockholder's equity before the stock dividend=23000000+33500000-10000000=$ 46500000
Total paid-in capital after stock dividend:
$
Common stock (650000*20)+240000 13240000
Preferred stock (100000*30) 3000000
Paid-in capital in excess of par-common stock (7000000+72000) 7072000
Paid-in capital in excess of par-preferred stock 0
Total paid-in capital 23312000
Total retained earnings after the stock dividend=33500000-312000=$ 33188000
Total stockholder's equity before the stock dividend=Total paid-in capital+Total retained earnings-Treasury stock=23312000+33188000-10000000=$ 46500000